Kellers Music was organized in January of 2017. The company offers music lessons to people of all ages and skill level. The company charges an annual $50 registration fee and $100 monthly due per student for music lessons. You have been hired by the company to install and maintain the company’s accounting system. The owner of Kellers Music would like see the following statements at the end of each month:

Income Statement
Balance Sheet
Retained Earnings Statement

The following transactions occurred during the month of January:

Jan. 1   Kellers Music is organized as a corporation and sells $50,000 of capital stock to its owners.
1 Paid the annual insurance premium of $1,824.
(Hint: Prepaids are recorded as an asset until the resource has been used up.)
1 Rented a building for $1,250 per month and paid the first three month’s rent in advance.
1   Paid $17,340 for the purchase, delivery, and installation of the following:

       Musical Instruments . . . . . . . . . . $12,000
       Computer . . . . . . . . . . . . . . . . . . . $1,440
       Furniture . . . . . . . . . . . . . . . . . . . $3,900
      
2 Received cash of $1,650 for signing up 11 students. The $1,650 included $550 collected for the
   annual registration fee and $1,100 for the monthly dues for lessons.
3 Paid $220 for fliers to be distributed this week to local schools.   
5    Purchased supplies (business cards, ink, paper, pencils, etc…) on account, $225.
5 Purchased snacks on account for grand opening event for $145 (advertising expense).
9 Received cash of $1,250 for signing up 10 additional students. The $1,250 included $500
collected for the annual registration fee and $750 for the prorated monthly dues for lessons.
10   Hired a music teacher to help out in the evenings at $20/hour. (Hint: No entry required.)
Signed up 7 additional students for $700. The $700 included $350 collected for the annual
registration fee and $350 for prorated monthly dues for lessons.
24 Paid the music teacher hired on January 10th for 9 hours.
26 Signed up 10 additional students which paid an annual registration fee of $50/student and
   $100/monthly dues. These students will not begin music lessons until next month, but each
student paid for February lessons in advance.
(Hint: The amount collected for the annual registration fee is earned in January as that is when
the registration took place in January. The amount collected for lessons is recorded as unearned
revenue since the service will not be provided until February).
31 The owner/manager is a musician giving lessons (employee of company) and takes a salary of
   $2,500 for first month
Answer #1
Income Statement
Particulars $ $
Registration Fees 1900 =550+500+350+(50*10)
Monthly Fees 2200 4100 =1100+750+350
Expenses:
Rent Expense 1250 =1250*1
Advertisement Expense 365 =220+145
Salary Expense 2680 4295 =(9*20)+2500
Net Income/(Loss) -195
Balance Sheet
Assets $ $
Non-Current Assets:
Musical Instruments 12000
Computer 1440
Furniture 3900 17340
Current Assets:
Cash 29141 =50000-1824-(1250*3)-17340+1650-220-145+1250+700-(9*20)+(50*10)+(100*10)-2500
Prepaid Insurance 1824
Prepaid Rent 2500 =1250*2
Supplies 225 33690
Total 51030
Equity & Liabilities $ $
Equity:
Common Stock 50000
Retained Earnings -195 49805
Liabilities:
Accounts Payable 225
Unearned Revenue 1000 1225
Total 51030
Retained Earnings Statement
Particulars $
Opening Balance 0
Net Income for the month -195
Closing Balance -195